# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v2

@@ 4.2.1. Integrating General Values to AI: how to incorporate general human values into AI development?
Instruction Data. This dimension refers to the types of data and processes used to provide guidance or direction to AI systems during their development. Human Annotation makes data with human-generated labels or annotations that indicate the presence or relevance of specific human values (Shen et al., 2023b; Pei et al., 2022). Human-AI CoAnnotation leverages both human expertise and AI capabilities to collaboratively annotate the data (Sun et al., 2023; Li et al., 2023d). Simulated Human Data generates synthetic or simulated data that mimics human behaviors, preferences, or decision-making processes to provide training signals for AI systems (Dubois et al., 2023; Kim et al., 2023a).
Model Learning. This dimension refers to the model architecture design and training stages after the data collection, where human values are integrated during the model learning process. Online Human Alignment integrates human values into AI systems in real-time or during active system operation, often through interactive feedback loops or adaptive learning mechanisms. Examples include real-time user feedback and online training (Ouyang et al., 2022; Dong et al., 2023b). Offline Human Alignment incorporates human values into AI systems prior to deployment or during offline training phases, without direct user interaction (Rafailov et al., 2023; Song et al., 2024; Yuan et al., 2023).
Inference Stage. This dimension involves evaluating the alignment of AI systems with human values and assessing their performance and behavior in relation to predefined criteria or benchmarks. Prompting leverages prompting methods, such as in-context learning and chain-of-thought, on trained AI systems to elicit or critique AI regarding the encoded values (Bai et al., 2022b). External Tool Interaction integrates external tools, like code interpreter for debugging, to cross-check and refine their initial generated content (Gou et al., 2024). Response Search generates a diverse range of high-quality outputs from which to choose (Bradley et al., 2024).
@@ 4.2.2. Customizing AI Values: how to customize AI to incorporate values from individuals or human groups?
Adapt Model by Learning. This dimension involves refining or adjusting AI values through techniques for customization, such as iterative learning process, model enhancements, or structural modifications. Group-based Learning trains AI models from specific user groups or communities to capture group-specific values or preferences (Zhao et al., 2023a). Active Learning aims to interative selecting and lebeling data samples for AI model training based on their potential to improve alignment with user preference or values (Peschl et al., 2022). Inserted Adapters incorporates adapter modules or components into AI model architectures to fine-tune specific aspects or behaviors (Sun et al., 2022). Mixture of Experts combines multiple specialized models or experts to collectively capture diverse perspectives and values, with each expert focusing on a specific subset of the data or problem space (Shen et al., 2022b). Enhanced Knowledge enhances AI model’s representations and embeddings with additional knowledge or context to improve alignment with user preferences or values (Wang et al., 2023c; Dalvi et al., 2022).
Interactive Alignment This dimension involves actively engaging users or stakeholders in the process of customizing AI values to align with specific contexts, domains, or user preferences. Interactive Learning enables the users to provide feedback or corrections to AI models in real time, such as using interactive tutorials and user-driven customization interfaces (Maghakian et al., 2023; Swazinna et al., 2023). Steering Prompts provides users with prompts or cues to steer the behavior or decision-making of AI systems towards desired outcomes or values (Lahoti et al., 2023; Fei et al., 2023). Proactive Alignment anticipates user needs and preferences based on historical data or user profiles and proactively adjusting AI systems accordingly (Shen et al., 2023a).
@@ 4.2.3. Evaluating AI Systems: how to evaluate AI regarding human values?
Human-In-The-Loop Evaluation. This dimension involves incorporating human judgement, feedback, or interaction into the evaluation process to assess the effectiveness, robustness, and ethical implications of integrating human values into AI systems. Human Evaluation solicits feedback, opinions, or assessments from human evaluators to gauge the alignment of AI systems with human values and ethical standards (Sharma et al., 2023; Ouyang et al., 2023; Kiesel et al., 2022; Hosking et al., 2024). Human-AI Collaborative Evaluation collaboratively evaluates AI systems with both human evaluators and large language models (LLMs) to leverage the strengths of both human judgement and AI capabilities (Kwon and Mihindukulasooriya, 2023a; Ammanabrolu et al., 2022).
Automatic Evaluation. This dimension involves using computational methods or algorithms to assess the alignment of AI systems with human values, without direct human involvement. Human Simulators use simulation models or virtual agents to mimic human behavior and assess AI performance in human-like scenarios. Typical methods include agent-based simulations, synthetic user models, and others (Jia et al., 2024; Wang et al., 2023d; Felkner et al., 2023). Evaluation Benchmarks establishes standardized benchmarks or metrics for evaluating AI performance in relation to human values and ethical considerations (Ramezani and Xu, 2023; Hu et al., 2024a; Kiesel et al., 2022; Rao et al., 2023; Lee et al., 2023a; Sheng et al., 2021; Sharma et al., 2023). Distribution Difference compares the difference between the output distribution from AI generations and human data to evaluate AI (Patel and Pavlick, 2021; Lee et al., 2023a).
@@ 4.2.4. Ecosystem and Platforms: how to build the ecosystem to facilitate human-AI alignment?
The ecosystem and platforms refer to the broader context in which AI systems operate and interact with other agents, platforms, or environments. This includes the infrastructure, frameworks, and technologies that support the development, deployment, and utilization of AI systems.
LLM-based Agents are based on large language models (LLMs) such as GPT (Generative Pre-trained Transformer) models, which have been pre-trained on vast amounts of text data (Dubois et al., 2023; Xu et al., 2023a; Wu et al., 2023a; Zhou et al., 2023b).
RL-based Agents are based on reinforcement learning (RL) algorithms to learn and adapt their behavior based on feedback from the environment or human users (Yuan et al., 2024).
Annotation Platforms refers to the ecosystems that are designed to crowdsource human demonstrations as collected data for reinforcement learning (Gerstgrasser et al., 2021) and supervised finetuning learning for alignment (Pei et al., 2022).
@@
ALIGN HUMANS to AI
When interacting with people, AI agents do not just influence the state of the world – they also influence the actions people take in response to the agent, and even their underlying intentions and strategies (Hong et al., 2023). From a long-term perspective, it is essential to consider the dynamic changes around human-AI alignment. Therefore, research in this direction aims to help humans better understand, critique, collaborate with, and adapt to AI advancements. The two core research questions we identified in this direction are: RQ3. Human Cognitive Adjustment to AI (Section 4.3) and RQ4. Human Adaptive Behavior to AI (Section 4.4).
4.3. Align Humans to AI:
@@ 4.3.1. Perceiving and Understanding AI: how do humans learn to perceive and explain AI systems?
Addressing the problem of human perception and understanding of AI includes the fundamental education and training required for less technical people to improve understanding of the behind mechanisms and outputs produced by AI systems. It also includes visualizations and human-centered explanations designed to help more people learn to understand AI outputs.
Educating and Training Humans. AI is increasingly incorporated into the daily lives of a broad spectrum of users, including those with little to no technical knowledge on how AI operates. Here we discuss how to help these less technical people become more AI literate, or better trained to use AI. AI Literacy and Awareness is broadly defined as the core competencies required for less technical people to better use and collaborate with AI (Long and Magerko, 2020). Beyond these core competencies, there is also more in depth and explicit Training Courses to support individuals to better collaborate and utilize AI (Paranjape et al., 2019).
AI Sensemaking and Explanations. To improve people’s understanding of AI systems and generated outputs, various techniques and approaches have been developed to help people learn to make sense of and explain the model’s outputs. Perception and Sensemaking on AI involves the process through which humans learn to make better sense of AI mechanisms and decision makings (Kaur et al., 2022; Sivaraman et al., 2023). People also need to understand how AI systems arrive at specific generated outputs. Prior studies in Human-Centered Explanations have examined various approaches and interactive techniques to increase human understanding of AI generations and outputs (Tintarev and Masthoff, 2007; Smith-Renner et al., 2020a).
@@ 4.3.2. Critical Thinking around AI: how do humans think critically about AI systems?
Trust and Reliance on AI Decisions refers to the extent to which people mentally trust the AI competency and practically rely on AI for output generation and decision making. We use the term Trustworthiness to indicate whether humans decide to trust the reliability, integrity, and competence of an AI system in delivering accurate and trustworthy decisions or recommendations to the users (Ma et al., 2023a; Srivastava et al., 2023). We define AI Reliability as to what extent humans utilize and rely on AI to automate decision making in practice with low error rates and robustness performance. (Schemmer et al., 2023; Vasconcelos et al., 2023; Suresh et al., 2022). Selective AI Adoption refers to the criteria and considerations that guide the adoption or rejection of AI technologies based on their perceived benefits, risks, and alignment with user needs and preferences (Gu et al., 2024; Vasconcelos et al., 2023).
@@
Re-Calibrating Cognition to Align with AI. This sub-category deals with interventions and techniques to help users adjust (1) their own mental model of how the AI operates and (2) recalibrate their perception of the AI. Updating Human Mental Model of AI means that when the AI is not aligned with humans, it is important to adjust human perceptions of the confidence in AI systems based on the model’s capability and track record (He et al., 2023c) . Recalibrating Trust and Reliability indicates that humans adjust their perceptions of trust and reliability in AI systems based on their performance and reliability. This is also important to foster appropriate reliance and skepticism Wischnewski et al. (2023).
4.4. Align Humans to AI:
@@
AI Assistant for Humans captures the essence of a symbiotic relationship where AI systems are designed to bolster human capabilities, with humans steering the interactions. LLM-based Demand Specification employs LLMs to interpret and respond to human requests, powering virtual assistants and chatbots that streamline information retrieval and improve task accuracy (Terry et al., 2023). This naturally extends to LLM-based Prompt Strategy, which helps humans to better write prompts using additional abstraction and scaffolding methods. The example system can enhance humans’ capability in generating intelligent prompts and suggestions, such as autocomplete and question-generation tools for LLM-based systems (Wu et al., 2022a; Wang et al., 2024), facilitating humans’ smooth and intuitive decision-making processes. In the creative arena, LLM-based Prototypes utilize AI to transform human ideas into tangible or organized concepts (Gebreegziabher et al., 2023), enabling professionals to explore and refine a wide array of artistic possibilities with AI-generated options (Huh et al., 2023). On a broader note, researchers have also explored how AI-Assisted Task and Decisions aims to achieve complementary performance for human-AI collaborative tasks by empower humans to discern when and how to adopt AI-assisted recommendations or AI-generated explanations for decision-making (Wang et al., 2022a; Wang and Yin, 2023; Zheng et al., 2023b).
Human-AI Partnership refers to a collaborative relationship where humans and AI systems work together as partners, combining their respective strengths to achieve shared goals more effectively. Simulated Agency enables humans to collaborate with AI partners with simulated agency or autonomy (e.g., autonomous agents and collaborative robots (Park et al., 2023)) to make decisions collaboratively. Reciprocal Learning focuses on how humans learn from and exchange knowledge with AI systems, enhancing human-AI collective capabilities and performance through knowledge-sharing platforms and collaborative filtering (Zagalsky et al., 2021). AI Delegation allows humans to delegate AI partners to help finish tasks or responsibilities, facilitated by task assignment algorithms and workflow management systems (Pinski et al., 2023). Furthermore, AI-Mediated Tasks emphasizes how traditional human tasks or behaviors (e.g., communication) would be changed by incorporating AI partners in the loop (Bernstein et al., 2010; Sundar and Lee, 2022; Xu and Zhang, 2023). Co-Design with AI explores AI as a design material or a partner in prototyping, which enables humans to converse with AI in situations that can collaboratively improve the design outcomes. (Yildirim et al., 2022; Wang et al., 2024)
@@ 4.4.2. AI Impact on Humans and Society: how are humans influenced by AI systems ?
Impacts on Participatory Individuals and Groups covers the effects of AI advancement on the behaviors, attitudes, and experiences of both individuals and groups. This dimension focuses on examining how AI technologies influence decision-making, creativity, privacy, and authorship. Decision Making refers to analyzing how AI technologies influence human decision-making processes, including biases, preferences, and risk assessment, in various domains such as healthcare, finance, and personal life (Shen and Huang, 2020). Human Creativity explores the impact of AI technologies on human creativity, innovation, and expression, including the augmentation or automation of creative tasks and the emergence of new forms of artistic expression (Ashkinaze et al., 2024). Privacy relates to investigating the implications of AI technologies for individual privacy rights, data protection, and surveillance, including concerns about data collection, tracking, and algorithmic profiling (Li et al., 2024b). Authorship catalogs issues related to intellectual property, attribution, and ownership of AI-generated content, including questions of legal responsibility, copyright infringement, and plagiarism detection. Salient is that AI can produce increasingly realistic, synthetic data quickly and at low cost, which brings forth tensions around the use of such data to make decisions (Hämäläinen et al., 2023).
Societal Concerns and AI Impacts involves the broader societal implications and consequences of AI advancement on misinformation, education, social relationships, norms, job displacement, and other aspects of human society. Misinformation and Moderation concerns the challenges of misinformation, disinformation, and online content moderation in the context of AI-driven information ecosystems, including concerns about algorithmic bias and filter bubbles (Atreja et al., 2023). Impacts on Education pertains to assessing the effects of AI technologies on education systems, learning outcomes, pedagogical practices, and workforce training, including opportunities for personalized learning and skill development (Kazemitabaar et al., 2023). Impacts on Social Relationship and Norms explores how AI technologies shape interpersonal relationships, social interactions, and cultural norms, including changes in communication patterns, social dynamics, and ethical considerations (Cheng et al., 2022b). Workplace refers to examining the effects of automation and AI technologies on employment patterns, job markets, and workforce dynamics, including concerns about job displacement, re-skilling, and economic inequality (Park et al., 2021a).
@@
Evaluate Human-AI Collaboration refers to the evaluation of the effectiveness of an AI system in collaboration with humans. It is key to not only consider the final output, but also the interaction experience (Lee et al., 2023d). Human-AI Team Performance compares human-AI team performance with the performance of humans alone without AI collaboration The metrics for measuring performance should include both task success metrics (e.g., accuracy) as well as indicators of efficiency (Mozannar et al., 2024; Ding et al., 2023a). Cognitive Workload and User Satisfaction involves understanding the degree of cognitive load that the user experiences when interacting with the system, as well as user satisfaction with the interaction and the final outcome. Such aspects are often captured via surveys or interviews (Ding et al., 2023a; Lee et al., 2023d). Control and Trust refers to how user control can support the avoidance of catastrophic AI failures, especially in high-stakes settings where AI mistakes could lead to harm (Ding et al., 2023b; Srikanth et al., 2024). Interaction Behavioral Analytics refers to measuring task performance quantitatively. This approach includes recording user interaction data and analyzing the patterns (Perry et al., 2023; Lee et al., 2022; Villalba et al., 2023). Qualitative Survey and Interviews refers to qualitative approaches to understanding human-AI interaction. Commonly used methods include qualitative survey questions (i.e., open-ended) and user interviews to assess aspects of the user experience. (Zhang et al., 2023c). Statistical Analysis utilizes methods such as regression analysis to quantitatively analyze and evaluate data from human studies, allowing for the verification of hypotheses (Wang et al., 2021).
@@
Evaluate Societal Impact refers to the macro-impact of a group of people as they come to use AI broadly. This dimension requires sufficient scale and time. The aim is to understand how the group’s behavior changes as people within it frequently interact with AI. Public Opinion Surveys aims to investigate the impact of AI on human measures of interest though deploying and analyzing large scale questionnaires (Santurkar et al., 2023). Behavioral Data Analytics collects large-scale and potentially longitudinal behavior data, with the aim of understanding how pattern evolve and shift over time (Shin et al., 2023; Zhou and Lee, 2024).
5. Paper Coding Analyses and Findings

### v3

@@ 4.2.1. Integrating General Values to AI: how to incorporate general human values into AI development?
Instruction Data. This dimension refers to the types of data and processes used to provide guidance or direction to AI systems during their development. Human Annotation makes data with human-generated labels or annotations that indicate the presence or relevance of specific human values (Shen et al., 2023b; Pei et al., 2022). Human-AI CoAnnotation leverages both human expertise and AI capabilities to collaboratively annotate the data (Sun et al., 2023; Li et al., 2023b). Simulated Human Data generates synthetic or simulated data that mimics human behaviors, preferences, or decision-making processes to provide training signals for AI systems (Dubois et al., 2023; Kim et al., 2023).
Model Learning. This dimension refers to the model architecture design and training stages after the data collection, where human values are integrated during the model learning process. Online Human Alignment integrates human values into AI systems in real-time or during active system operation, often through interactive feedback loops or adaptive learning mechanisms. Examples include real-time user feedback and online training (Ouyang et al., 2022; Dong et al., 2023). Offline Human Alignment incorporates human values into AI systems prior to deployment or during offline training phases, without direct user interaction (Rafailov et al., 2023; Song et al., 2024; Yuan et al., 2023).
Inference Stage. This dimension involves evaluating the alignment of AI systems with human values and assessing their performance and behavior in relation to predefined criteria or benchmarks. Prompting leverages prompting methods, such as in-context learning and chain-of-thought, on trained AI systems to elicit or critique AI regarding the encoded values (Bai et al., 2022b; Wang et al., 2023a). External Tool Interaction integrates external tools, like code interpreter for debugging, to cross-check and refine their initial generated content (Gou et al., 2024). Response Search generates a diverse range of high-quality outputs from which to choose (Bradley et al., 2024).
@@ 4.2.2. Customizing AI Values: how to customize AI to incorporate values from individuals or human groups?
Adapt Model by Learning. This dimension involves refining or adjusting AI values through techniques for customization, such as iterative learning process, model enhancements, or structural modifications. Group-based Learning trains AI models from specific user groups or communities to capture group-specific values or preferences (Zhao et al., 2023). Active Learning aims to interative selecting and lebeling data samples for AI model training based on their potential to improve alignment with user preference or values (Peschl et al., 2022). Inserted Adapters incorporates adapter modules or components into AI model architectures to fine-tune specific aspects or behaviors (Sun et al., 2022). Mixture of Experts combines multiple specialized models or experts to collectively capture diverse perspectives and values, with each expert focusing on a specific subset of the data or problem space (Shen et al., 2022). Enhanced Knowledge enhances AI model’s representations and embeddings with additional knowledge or context to improve alignment with user preferences or values (Wang et al., 2023c; Dalvi et al., 2022).
Interactive Alignment This dimension involves actively engaging users or stakeholders in the process of customizing AI values to align with specific contexts, domains, or user preferences. Interactive Learning enables the users to provide feedback or corrections to AI models in real time, such as using interactive tutorials and user-driven customization interfaces (Maghakian et al., 2023; Swazinna et al., 2023). Steering Prompts provides users with prompts or cues to steer the behavior or decision-making of AI systems towards desired outcomes or values (Lahoti et al., 2023; Fei et al., 2023). Proactive Alignment anticipates user needs and preferences based on historical data or user profiles and proactively adjusting AI systems accordingly (Welch et al., 2022; Shen et al., 2023a).
@@ 4.2.3. Evaluating AI Systems: how to evaluate AI regarding human values?
Human-In-The-Loop Evaluation. This dimension involves incorporating human judgement, feedback, or interaction into the evaluation process to assess the effectiveness, robustness, and ethical implications of integrating human values into AI systems. Human Evaluation solicits feedback, opinions, or assessments from human evaluators to gauge the alignment of AI systems with human values and ethical standards (Liu et al., 2022; Sharma et al., 2023; Ouyang et al., 2023; Kiesel et al., 2022; Hosking et al., 2024). Human-AI Collaborative Evaluation collaboratively evaluates AI systems with both human evaluators and large language models (LLMs) to leverage the strengths of both human judgement and AI capabilities (Kwon and Mihindukulasooriya, 2023; Ammanabrolu et al., 2022).
Automatic Evaluation. This dimension involves using computational methods or algorithms to assess the alignment of AI systems with human values, without direct human involvement. Human Simulators use simulation models or virtual agents to mimic human behavior and assess AI performance in human-like scenarios. Typical methods include agent-based simulations, synthetic user models, and others (Jia et al., 2024; Wang et al., 2023d; Felkner et al., 2023; Ruan et al., 2023). Evaluation Benchmarks establishes standardized benchmarks or metrics for evaluating AI performance in relation to human values and ethical considerations (Sun et al., 2023; Ramezani and Xu, 2023; Hu et al., 2024; Kiesel et al., 2022; Rao et al., 2023; Lee et al., 2023a; Sheng et al., 2021; Sharma et al., 2023). Distribution Difference compares the difference between the output distribution from AI generations and human data to evaluate AI (Patel and Pavlick, 2021; Lee et al., 2023a).
@@ 4.2.4. Ecosystem and Platforms: how to build the ecosystem to facilitate human-AI alignment?
The ecosystem and platforms refer to the broader context in which AI systems operate and interact with other agents, platforms, or environments. This includes the infrastructure, frameworks, and technologies that support the development, deployment, and utilization of AI systems. LLM-based Agents are based on large language models (LLMs) such as GPT (Generative Pre-trained Transformer) models, which have been pre-trained on vast amounts of text data (Dubois et al., 2023; Xu et al., 2023; Wu et al., 2023a; Zhou et al., 2023b). RL-based Agents are based on reinforcement learning (RL) algorithms to learn and adapt their behavior based on feedback from the environment or human users (Yuan et al., 2024). Annotation Platforms refers to the ecosystems that are designed to crowdsource human demonstrations as collected data for reinforcement learning (Gerstgrasser et al., 2021) and supervised finetuning learning for alignment (Pei et al., 2022).
| Underexplored Dimensions in Aligning AI to Humans (See Sec 5.1 for supporting data and evidence): 1. Implicit human feedback and simulated human value feedback are under-explored in existing research work. 2. Developing and customizing AI during the inference stage or in an interactive way is under-explored. 3. Human-in-the-loop evaluation is much less explored than automatic evaluation. |
@@
ALIGN HUMANS to
AI
When interacting with people, AI agents do not just influence the state of the world – they also influence the actions people take in response to the agent, and even their underlying intentions and strategies (Hong et al., 2023). From a long-term perspective, it is essential to consider the dynamic changes around human-AI alignment. This direction outlines alignment research from the Human-centered perspective (e.g., HCI/Social Science domains) and provides HCI researchers and user experience designers with guidance for approaching two core research questions: RQ3. Human Cognitive Adjustment to AI (Section 4.3) and RQ4. Human Adaptive Behavior to AI (Section 4.4).
4.3. Align Humans to
AI:
@@ 4.3.1. Perceiving and Understanding AI: how do humans learn to perceive and explain AI systems?
Addressing the problem of human perception and understanding of AI includes the fundamental education and training required for less technical people to improve understanding of the behind mechanisms and outputs produced by AI systems. It includes visualizations and human-centered explanations tohelp more people learn to understand AI outputs.
Educating and Training Humans. AI is increasingly incorporated into the daily lives of a broad spectrum of users, including those with little to no technical knowledge on how AI operates. Here we discuss how to help these less technical people become more AI literate, or better trained to use AI. AI Literacy and Awareness is broadly defined as the core competencies required for less technical people to better use and collaborate with AI (Long and Magerko, 2020). Beyond, there is also more in depth and explicit Training Courses to support individuals to better collaborate and utilize AI (Paranjape et al., 2019; McDonald and Pan, 2020).
AI Sensemaking and Explanations. To improve people’s understanding of AI systems and generated outputs, various techniques and approaches have been developed to help people learn to make sense of and explain the model’s outputs. Perception and Sensemaking on AI involves the process through which humans learn to make better sense of AI mechanisms and decision makings (Kaur et al., 2022; Sivaraman et al., 2023; Petridis et al., 2023). People also need to understand how AI systems arrive at specific generated outputs. Prior studies in Human-Centered Explanations have examined various approaches and interactive techniques to increase human understanding of AI generations and outputs (Tenney et al., 2020; Tintarev and Masthoff, 2007; Smith-Renner et al., 2020).
@@ 4.3.2. Critical Thinking around AI: how do humans think critically about AI systems?
Trust and Reliance on AI Decisions refers to the extent to which people mentally trust the AI competency and practically rely on AI for output generation and decision making. We use the term Trustworthiness to indicate whether humans decide to trust the reliability, integrity, and competence of an AI system in delivering accurate and trustworthy decisions or recommendations to the users (Yin et al., 2019; Ma et al., 2023a; Srivastava et al., 2023). We define AI Reliability as to what extent humans utilize and rely on AI to automate decision making in practice with low error rates and robustness performance (Schemmer et al., 2023). (Schemmer et al., 2023; Vasconcelos et al., 2023; Suresh et al., 2022). Selective AI Adoption refers to the criteria and considerations that guide the adoption or rejection of AI technologies based on their perceived benefits, risks, and alignment with user needs and preferences (Gu et al., 2024; Vasconcelos et al., 2023).
@@
Re-Calibrating Cognition to Align with AI. This sub-category deals with interventions and techniques to help users adjust (1) their own mental model of how the AI operates and (2) recalibrate their perception of the AI. Updating Human Mental Model of AI means that when the AI is not aligned with humans, it is important to adjust human perceptions of the confidence in AI systems based on the model’s capability and track record (He et al., 2023; Kocielnik et al., 2019) . Recalibrating Trust and Reliability indicates that humans adjust their perceptions of trust and reliability in AI systems based on their performance and reliability. This is also important to foster appropriate reliance and skepticism (Wischnewski et al., 2023).
4.4. Align Humans to
AI:
@@ 4.4. Align Humans to
AI Assistant for Humans captures the essence of a symbiotic relationship where AI systems are designed to bolster human capabilities, with humans steering the interactions. LLM-based Demand Specification employs LLMs to interpret and respond to human requests, powering virtual assistants and chatbots that streamline information retrieval and improve task accuracy (Terry et al., 2023). This naturally extends to LLM-based Prompt Strategy, which helps humans to better write prompts using additional abstraction and scaffolding methods. The example system can enhance humans’ capability in generating intelligent prompts and suggestions, such as autocomplete and question-generation tools for LLM-based systems (Wu et al., 2022; Wang et al., 2024), facilitating humans’ smooth and intuitive decision-making processes. In the creative arena, LLM-based Prototypes utilize AI to transform human ideas into tangible or organized concepts (Gebreegziabher et al., 2023), enabling professionals to explore and refine a wide array of artistic possibilities with AI-generated options (Huh et al., 2023). On a broader note, researchers have also explored how AI-Assisted Task and Decisions aims to achieve complementary performance for human-AI collaborative tasks by empower humans to discern when and how to adopt AI-assisted recommendations or AI-generated explanations for decision-making (Choi et al., 2019; Wang et al., 2022; Wang and Yin, 2023; Zheng et al., 2023; Wu et al., 2023c).
Human-AI Partnership refers to a collaborative relationship where humans and AI systems work together as partners, combining their respective strengths to achieve shared goals more effectively. Simulated Agency enables humans to collaborate with AI partners with simulated agency or autonomy (e.g., autonomous agents and collaborative robots (Park et al., 2023)) to make decisions collaboratively. Reciprocal Learning focuses on how humans learn from and exchange knowledge with AI systems, enhancing human-AI collective capabilities and performance through knowledge-sharing platforms and collaborative filtering (Zagalsky et al., 2021; Pinski et al., 2023). AI Delegation allows humans to delegate AI partners to help finish tasks or responsibilities, facilitated by task assignment algorithms and workflow management systems (Pinski et al., 2023). Furthermore, AI-Mediated Tasks emphasizes how traditional human tasks or behaviors (e.g., communication) would be changed by incorporating AI partners in the loop (Bernstein et al., 2010; Sundar and Lee, 2022; Xu and Zhang, 2023). Co-Design with AI explores AI as a design material or a partner in prototyping, which enables humans to converse with AI in situations that can collaboratively improve the design outcomes. (Yildirim et al., 2022; Wang et al., 2024)
@@ 4.4.2. AI Impact on Humans and Society: how are humans influenced by AI systems ?
Impacts on Participatory Individuals and Groups covers the effects of AI advancement on the behaviors, attitudes, and experiences of both individuals and groups. This dimension focuses on examining how AI technologies influence decision-making, creativity, privacy, and authorship. Decision Making refers to analyzing how AI technologies influence human decision-making processes, including biases, preferences, and risk assessment, in various domains such as healthcare, finance, and personal life (Shen and Huang, 2020). Human Creativity explores the impact of AI technologies on human creativity, innovation, and expression, including the augmentation or automation of creative tasks and the emergence of new forms of artistic expression (Ashkinaze et al., 2024). Privacy relates to investigating the implications of AI technologies for individual privacy rights, data protection, and surveillance, including concerns about data collection, tracking, and algorithmic profiling (Li et al., 2024a). Authorship catalogs issues related to intellectual property, attribution, and ownership of AI-generated content, including questions of legal responsibility, copyright infringement, and plagiarism detection. Salient is that AI can produce increasingly realistic, synthetic data quickly and at low cost, which brings forth tensions around the use of such data to make decisions (Hämäläinen et al., 2023; Atreja et al., 2023).
Societal Concerns and AI Impacts involves the broader societal implications and consequences of AI advancement on misinformation, education, social relationships, norms, job displacement, and other aspects of human society. Misinformation and Moderation concerns the challenges of misinformation, disinformation, and online content moderation in the context of AI-driven information ecosystems, including concerns about algorithmic bias and filter bubbles (Atreja et al., 2023). Impacts on Education pertains to assessing the effects of AI technologies on education systems, learning outcomes, pedagogical practices, and workforce training, including opportunities for personalized learning and skill development (Kazemitabaar et al., 2023). Impacts on Social Relationship and Norms explores how AI technologies shape interpersonal relationships, social interactions, and cultural norms, including changes in communication patterns, social dynamics, and ethical considerations (Cheng et al., 2022). Workplace refers to examining the effects of automation and AI technologies on employment patterns, job markets, and workforce dynamics, including concerns about job displacement, re-skilling, and economic inequality (Park et al., 2021a).
@@
Evaluate Human-AI Collaboration refers to the evaluation of the effectiveness of an AI system in collaboration with humans. It is key to not only consider the final output, but also the interaction experience (Lee et al., 2023b). Human-AI Team Performance compares human-AI team performance with the performance of humans alone without AI collaboration The metrics for measuring performance should include both task success metrics (e.g., accuracy) as well as indicators of efficiency (Mozannar et al., 2024; Ding et al., 2023a). Cognitive Workload and User Satisfaction involves understanding the degree of cognitive load that the user experiences when interacting with the system, as well as user satisfaction with the interaction and the final outcome. Such aspects are often captured via surveys or interviews (Ding et al., 2023a; Lee et al., 2023b). Control and Trust refers to how user control can support the avoidance of catastrophic AI failures, especially in high-stakes settings where AI mistakes could lead to harm (Srikanth et al., 2024; Ding et al., 2023b). Interaction Behavioral Analytics refers to measuring task performance quantitatively. This approach includes recording user interaction data and analyzing the patterns (Villalba et al., 2023; Perry et al., 2023). Qualitative Survey and Interviews refers to qualitative approaches to understanding human-AI interaction. Commonly used methods include qualitative survey questions (i.e., open-ended) and user interviews to assess aspects of the user experience. (Zhang et al., 2023b). Statistical Analysis utilizes methods such as regression analysis to quantitatively analyze and evaluate data from human studies, allowing for the verification of hypotheses (Wang et al., 2021).
@@
Evaluate Societal Impact refers to the macro-impact of a group of people as they come to use AI broadly. This dimension requires sufficient scale and time. The aim is to understand how the group’s behavior changes as people within it frequently interact with AI. Public Opinion Surveys aims to investigate the impact of AI on human measures of interest though deploying and analyzing large scale questionnaires (Santurkar et al., 2023). Behavioral Data Analytics collects large-scale and potentially longitudinal behavior data, with the aim of understanding how pattern evolve over time (Shin et al., 2023; Zhou and Lee, 2024).
| Underexplored Dimensions in Aligning Humans to AI (See Sec 5.1 for supporting data and evidence): 1. Educating and training humans on AI literacy is under-investigated. 2. Auditing AI for various ethical values is not fully explored. 3. The collaboration between humans and AI with similar or superior capabilities is under-explored. 4. The societal impacts of and reactions to AI advancements are not fully explored. |
5. Findings and Discussions on Current Gaps